OpenStudy (anonymous):

6 square 32 - 4 square root 2

jimthompson5910 (jim_thompson5910):

is the problem \[\Large 6\sqrt{32} - 4\sqrt{2}\] ???

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

\[\Large 6\sqrt{32} - 4\sqrt{2}\] \[\Large 6\sqrt{16*2} - 4\sqrt{2}\] \[\Large 6\sqrt{16}*\sqrt{2} - 4\sqrt{2}\] \[\Large 6*4*\sqrt{2} - 4\sqrt{2}\] \[\Large 24*\sqrt{2} - 4\sqrt{2}\] \[\Large (24 - 4)\sqrt{2}\] \[\Large 20\sqrt{2}\] ---------------------------------------------- So \[\Large 6\sqrt{32} - 4\sqrt{2} = 20\sqrt{2}\]

OpenStudy (anonymous):

oh ok ty now i understand it